OpenStudy (anonymous):

The polar coordinates of a point are given. Find the rectangular coordinates of each point. (5,(π / 4)) (-2,(π / 6))

OpenStudy (johnweldon1993):

Or instead of going the long way backwards...we can also know that \[\large x = rcos(\theta)\] and \[\large y = rsin(\theta)\]

OpenStudy (johnweldon1993):

So for example with your first point up there \[\large x = 5cos(\pi/4)\] \[\large y = 5sin(\pi/4)\]

OpenStudy (anonymous):

And I can do the same for the second equation @johnweldon1993

OpenStudy (johnweldon1993):

Correct indeed :)

OpenStudy (anonymous):

so \[x=-2(\cos(\pi/6)\]

OpenStudy (johnweldon1993):

I just deleted the first post I made so as not to confuse you...why go a long way when there are shortcuts! lol And yes that is correct for your x-coordinate!

OpenStudy (anonymous):

Thank you so much @johnweldon1993 !

OpenStudy (johnweldon1993):

Anytime!

OpenStudy (anonymous):

@Johnweldon1993 is it weird that when I go to convert the first they come out to be the same point?

OpenStudy (johnweldon1993):

Not weird at all :P Seeing as how \(\large sin(45) = cos(45)\) :)

OpenStudy (anonymous):

Ah, thank you! All this polar stuff is just alsfjasflkajlfa! @johnweldon1993

OpenStudy (johnweldon1993):

Lol it will come much easier dont worry :) Actually polar is the easiest of the bunch...wait till you get to cylindrical and spherical THOSE are the best ;) haha

OpenStudy (anonymous):

So I'm off to go dig my shallow grave.

OpenStudy (johnweldon1993):

Lol well if you are in DE right now, you wont deal with them any if at all...Calc 3 is where that comes into play :) So nope, stay above ground...for now :P lol no try and stay permanently lol :D